# 2D Entanglement from a Twisted Motzkin Spaghetto

### v2

@@ Abstract
While highly entangled ground states of gapless local Hamiltonians have been known to exist in one dimension, their 2D counterparts were only recently found, with rather sophisticated interactions involving at least 4 neighboring degrees of freedom. Here, we show that similar 2D entanglement properties can be realized by twisting a single 1D chain embedded on a 2D lattice. The interaction to generate such 2D entanglement is much simpler than the previous 2D construction by coupling orthogonal arrays of 1D highly entangled chains. The new construction exhibits an entanglement phase transition with enhanced entanglement entropy at the critical point compared to the previous bona fide 2D models, but the spectral gap in the highly entangled phase decays faster. The tensor network representation of the new ground state is composed of tensors with lower rank, showing a similar global geometry as those of the old ones.
I Introduction
@@
It is not so straightforward to generalize the physical picture of Bell pair from 1D to 2D, as entanglement in two different directions turns out to compete against each other. A naive generalization to the rainbow chain was made in Ref. [18], only to end up with a lattice of Hausdorff dimension one embedded in 2D space. The trade-off between entanglement in orthogonal directions was also exemplified in colored generalizations of the toric code model [19, 20], where enhanced entanglement along the bipartition boundary results in anomalous topological EE of the GS. Another obstacle is a well-defined notion of height in 2D, when height differences can be counted along different paths connecting two points in the lattice. The prescription to fix both difficulties is to couple two orthogonal arrays of 1D Motzkin or Fredkin chains with appropriate vertex or tiling rules, such that a Coulomb gauge emerges to define an unambiguous height function [21, 22]. Moreover, the coupling leaves the Hamiltonian frustration free, and do not interfere with the interactions along each entangled chain. So unlike the anisotropic coupled wire construction of topological superconductors [23] where the inter-wire coupling opens up a gap, the system remains gapless. The resulting GSs are superpositions of discrete random membranes with a hard wall constraint at zero height, weighted by the deformation parameter to the power of the volume under the surfaces. The same entanglement phase transition occurs, except that the reduced height fluctuation in higher dimensions results in a logarithmic violation to the area law at the critical point instead of power law as in the 1D models.
@@
While the six- or 19-vertex and lozenge tiling realizations appear to be the natural generalization of the highly entangled Motzkin (resp. Fredkin) chain to 2D, the support of the local Hamiltonian operators grows from two (resp. three) in 1D to four (resp. six). So while the paradigm in principle works for higher dimensions, the interactions become increasingly less realistic. There has also been a proposal of sequentially generable 2D coupled GS that is highly entangled in only one direction with local Hamiltonian operators having even larger support of eight [24]. An alternative loop-soup construction was proposed in Ref. [19], where the GS is a superposition of random loop configurations on a 2D lattice, and on each loop lives a periodic Motzkin chain [25]. However, due to the randomness of the loop configurations, it is not clear how the EE for any bipartition would actually scale with the subsystem size. Moreover, a naive generalization of the Motzkin Hamiltonian to periodic boundary conditions leaves the GS with exponential degeneracy, making a combination with exotic EE scaling trivially exist. To avoid both of those complications, we focus on a twisting open 1D chain from the center to fill a 2D lattice, such that entanglement in both directions of the 2D lattice is isotropic. At the same time, the interaction is turned on only for 2D lattice bonds that reside along the twisted chain, so it remains two or three nearest neighbors. Since the location of the entangling chain is fixed, we are able to argue from the 1D results that the EE scaling is asymptotically the same as the coupled-chain construction, except at the critical point.
@@
The rest of the article is organized as follows. Sec. II defines the model with highly entangled GS in two directions by filling a 2D square lattice with a Motzkin chain twisted from the center. Sec. III analyzes the asymptotic scaling of the EE by mapping bipartitions of the 2D system into disjoint segments along the 1D chain. The result together with the scaling behavior of spectral gap converted from the 1D results for different phases and the critical points are compared with those for the previous coupled-chain construction. Sec. IV highlights the induced antiferromagnetic order in the spin sector along the radial direction of the 2D lattice. Sec. V introduces a new TN representation of the highly entangled 1D chains, the 2D twisted version of which is compared with the genuine 2D TN with similar global structure. Finally, a conclusion is given in Sec. VI with outlook on higher dimensional generalizations and EE scaling laws for gapless GSs.
II Twisted Motzkin chain
The $q$-deformed bicolor Motzkin chain [15] has a five-dimensional local Hilbert space, with spin up and down, both in either red (r) or blue (b) color, and a colorless spin zero. The Hamiltonian is expressed in terms of projectors $P_{j}^{c}=|L^{c}\rangle_{j,j+1}\langle L^{c}|_{j,j+1}+|R^{c}\rangle_{j,j+1}\langle R^{c}|_{j,j+1}+|F^{c}\rangle_{j,j+1}\langle F^{c}|_{j,j+1}$, where
@@ II Twisted Motzkin chain
Since the 2D model is constructed by twisting a 1D chain quadratically as long, we only need to specify the bonds on the 2D lattice occupied by degrees of freedom along the 1D chain. This is shown in Fig. 1 (a). The colored spins reside only on the edges of the square lattice covered by the twisted 1D chain, and interactions are only turned on between adjacent spins on the chain. Depending on the location of the spin in the square lattice, it can be either between east-west (resp. north-south) neighbors for the upper and lower (resp. left and right) quadrants. Near the two diagonal lines, where the winding 1D chain turns, the interaction is between neighbors around a corner.
@@
Figure 1: (a) A twisted Motzkin chain of length $2N=L(L+2)$ embedded in an $L\times L$ square lattice, shown here for $L=8$. The spin degrees of freedoms live on the edges of the square lattice, as are numbered in the figure. Four symmetric bipartition cuts are shown in different colors. (b) The two subsystems resulting from the yellow cut in (a). The northeast subsystem is represented with the thickened black lines, whereas the southwest is marked by thin gray lines. The configuration depicted corresponds to the probabilistically dominant Motzkin path in the GS superposition for the $q>1$ phase, to illustrate the entanglement across subsystems. Typical height configuration at the critical point is depicted in blue.
@@
To study the isotropic EE, four symmetric bipartition cuts are shown in Fig. 1 (a). They separate the spins into two subsystems consisting of increasingly larger segments of size 4na4n-a, for $n=1,2,\cdots,L/2+1$, with $a=0,1,2,3$ corresponding respectively to the yellow, blue, green and red cut. Starting from the center, segments within one subsystem in the 2D lattice on a given side of the cut alternate between the first and second halves of the 1D chain when it is stretched straight. Obviously the scaling behavior of bipartite EE for any of the cuts will not differ qualitatively, so from now on we focus on the cut along the main diagonal with $a=0$, which results in a bipartition shown in Fig. 1 (b).
III Scaling of EE and spectral gap
@@
As analyzed in Ref. [15], the EE of the Motzkin GS decomposes into a leading contribution from the color degree of freedom, and a subleading contribution from the spin sector. Only up-down spin pairs belonging to different subsystems account for the former. In the Motzkin path picture, this means that only the net height change of a subsystem, or the surplus of up spins matters. For the 1D system studied there, the leading contribution to the EE is therefore proportional to the height at the center of the Motzkin path, giving the three distinct scaling behaviors in the two different phases and the critical point. For the critical point at $q=1$, the EE of a single connected segment in the middle of the chain has also been obtained from combinatorial analyses for the colorless [30] and colored Motzkin chains [31].
Here, the situation is a bit more complicated, as segments belonging to two subsystems alternate along the chain. But we can get lower or upper bounds for the two phases with the following argument. For $q>1$, the net height change within each segment scales linearly with its length, at least for segments long enough which are the overwhelming majority in the asymptotic limit. If there is a dip in the middle of the segment, it would only increase the entanglement between the two subsystems, as the immediate neighboring segment on the left belongs to the other subsystem. For down spins to actually form up-down pairs with up spins from the next neighboring segment to the left, coming from the same subsystem, the probability would be of the next subleading order. Therefore, assuming the height to monotonically increase for the first half of the chain gives a lower bound on the EE, which is already growing linearly following the proof for the 1D case [15].
For $q<1$, the net height fluctuation stays at order $\mathcal{O}(1)$ anywhere along the chain. In the best case scenario for accumulating EE, the height changes are positive for all the segments in one subsystem, and negative in the other. Since there are in total $L/2$ segments of each, the EE scales at most with the linear size of the lattice, which means the area law for the 2D system.
At $q=1$, the net height change for a segment of size $4n$ typically scales as $\sqrt{4n}$. This can be understood as the standard deviation of a random walk of $4n$ steps, and is qualitatively unaffected by the hard wall constraint from below. So the total number of up-down spins to be color coordinated across two subsystems can be estimated as $\sum_{n=1}^{L/2}\sqrt{4n}$, which scales as $L^{3/2}$. The difference from the $L\log L$ scaling of the couple chain constructions [21, 22] can be understood from the enhanced fluctuation of height in 1D interaction compared to 2D.
@@
The subleading contribution to the EE is from the fluctuation in the height of Motzkin paths, which is also the sole source of entanglement for a monochromatic version of the model. A Schmidt decomposition of the reduced density matrix shows that the two subsystems only depend on each other via the total spin within one subsystem, such that the total spin of the whole chain is zero. Moving from the 1D system to our 2D construction, nothing changes except the subsystem is distributed into disconnected segments. So dimensionality of the reduced density matrix is still the number of possible values of the total spin. However, in the $q>1$ (resp. $q<1$) phase, the probability of the total spin, or height change decays exponentially with the deviation of the area underneath a Motzkin path from the most probable one that is of highest height (resp. completely flat) [15, 16]. So only paths with a height change at finite distance away from the typical one effectively matters, given a bounded EE independent of system size for both phases. When $q=1$, there is more fluctuation. Although entropic repulsion pushes the average height to order $\sqrt{N}$ above the hard wall at zero height, the variance of the Gaussian free field remains to scale linearly with the total length of the segments, making the EE scale as $\log N$ [11]. Note that in a monochromatic version of the 2D model, all three scalings are sub-area law, which is unique to our construction of filling the 2D lattice with a twisted 1D chain. The EE scalings for the two different 2D models are compared in Table 1 for different phases.
@@
Table 1: Comparison of the EE scalings with linear system size $L$ between the two different 2D constructions.
Model Phase Coupled chains Twisted chain
$q>1$ $L^{2}$ [21, 22] $L^{2}$
Colored $q=1$ $L\log L$ [21, 22] $L^{3/2}$
$q<1$ $L$ [21, 22] $L$
$q>1$ $\mathcal{O}(1)$ $\mathcal{O}(1)$
@@
Colorless $q=1$ $L$  [^1] This is an upper-bound from the zeroth order Rényi entropy. The random membrane when bipartitioned gives a Motzkin path on the boundary, giving the dimensionality of the reduced density matrix. The same upper-bound was also obtained for six-vertex models with domain wall boundary conditions using the exact enumeration of alternating sign matrices [32]. $\log L$
$q<1$ $\mathcal{O}(1)$ $\mathcal{O}(1)$
Table 2: Comparison of upper-bounds on the scaling of spectral gap with the linear system size $L$.
Model Phase Coupled chains Twisted chain
$q>1$ $e^{-\alpha L^{3}}$ [21] $e^{-\beta L^{4}}$ [33, 17]
Colored $q=1$ $L^{-\gamma}$  [^1] Conjectured in Ref. [29]. $L^{-\delta}$ [12]
$q<1$ $\mathcal{O}(1)$ 11footnotemark: 1 $\mathcal{O}(1)$ [34]
$q>1$ $e^{-\epsilon L^{2}}$ [21] $e^{-\zeta L^{2}}$ [33, 17]
Colorless $q=1$ $L^{-\eta}$ 11footnotemark: 1 $L^{-\theta}$ [11]
$q<1$ $\mathcal{O}(1)$ 11footnotemark: 1 $\mathcal{O}(1)$  [34]
@@
Unlike the scaling of EE discussed above, which requires somewhat different analyses than the 1D system due to the more complicated bipartition, upper-bounds on the scaling of the spectral gap can be straightforwardly deduced from the 1D scaling simply by substituting the linear system size with $N=L(L+2)$. The results are summarized in Table 1. This should be compared with the EE scaling in Table 1. In particular, the $q>1$ phase of the colored model shows that the same EE scaling does not imply the same scaling of spectral gap, and vice versa as shown by the critical point of the colored model and the $q>1$ phase of the colorless one.
IV Antiferromagnetic order

### v3

# The Motzkin Spaghetto

@@ Abstract
While highly entangled ground states of gapless local Hamiltonians have been known to exist in one dimension, their two-dimensional counterparts were only recently found, with rather sophisticated interactions involving at least four neighboring degrees of freedom. Here, we show that similar bipartite entanglement properties can be realized on a square lattice with anisotropic interactions in four different quadrants. The interaction to generate such entanglement is much simpler than the previous constructions by coupling orthogonal arrays of highly entangled chains. The new construction exhibits an entanglement phase transition with different scalings of entanglement entropy at the critical point and in the lowly entangled phase, and faster decay of the spectral gap in the highly entangled phase. The tensor network representation of the new ground state consists of tensors with lower rank, while preserving a global geometry similar to that of the original networks.
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
@@
It is not so straightforward to generalize the physical picture of Bell pair from 1D to 2D, as entanglement in two different directions competes against each other due to the monogamy of entanglement. Naive generalizations to the rainbow chain have been made in Ref. [10] and [18], which either exhibits volume scaling entanglement in only one of the two directions on a square lattice, or along in both directions of a lattice of Hausdorff dimension one embedded in 2D space. The trade-off between entanglement in orthogonal directions was also exemplified in colored generalizations of the toric code model [19, 20], where enhanced entanglement along the bipartition boundary results in anomalous topological EE of the GS.
@@
Another obstacle is a well-defined notion of height in 2D, when height differences can be counted along different paths connecting two points in the lattice. The prescription to fix both difficulties is to couple two orthogonal arrays of 1D Motzkin or Fredkin chains with appropriate vertex or tiling rules, such that a Coulomb gauge emerges to define an unambiguous height function [21, 22]. Moreover, the coupling leaves the Hamiltonian frustration free, and do not interfere with the interactions along each entangled chain. So unlike the anisotropic coupled wire construction of topological superconductors [23] where the inter-wire coupling opens up a gap, the system remains gapless. The resulting GSs are superpositions of discrete random membranes with a hard wall constraint at zero height, weighted by the deformation parameter to the power of the volume under the surfaces. The same entanglement phase transition occurs, except that the reduced height fluctuation in higher dimensions results in a logarithmic violation to the area law at the critical point instead of power law as in the 1D models.
@@
While the six- or 19-vertex and lozenge tiling realizations appear to be the natural generalization of the highly entangled Motzkin (resp. Fredkin) chain to 2D, the support of the local Hamiltonian operators grows from two (resp. three) in 1D to four (resp. six). So while the paradigm in principle works for higher dimensions, the interactions become increasingly less realistic. There has also been a proposal of sequentially generable 2D coupled GS that is highly entangled in only one direction with local Hamiltonian operators having even larger support of eight [24]. An alternative loop-soup construction was proposed in Ref. [19], where the GS is a superposition of random loop configurations on a 2D lattice, and on each loop lives a periodic Motzkin chain [25]. However, due to the randomness of the loop configurations, it is not clear how the EE for any bipartition would actually scale with the subsystem size. Moreover, a naive generalization of the Motzkin Hamiltonian to periodic boundary conditions leaves the GS with exponential degeneracy [26], making a combination with exotic EE scaling trivially exist.
Both of those complications turn out to be absent with another route towards a well-defined height variable in 2D, namely by specifying a fixed path of counting height increments. For a traversal of the 2D plane along a path twisted from its center, this amounts to introducing anisotropic interactions on a square lattice along different directions in different quadrants. The choice of the neighboring degrees of freedom is such that their bond traces out a Motzkin chain twisted from its center. Since the location of the entangling degrees of freedom is fixed, the EE scaling property is analytically tractable from the asymptotic EE scaling behavior of segments of the Motzkin chain with lengths forming an arithmetic series. Like the EE scaling of the coupled-chain models, the bipartite EE of the current construction in the highly entangled phase scales with the volume of the subsystem regardless of the direction of the bipartition cut. But unlike the previous models, here EE obeys the area law at the critical point, and does not grow with the subsystem size in the lowly entangled phase.
@@
The rest of the article is organized as follows. Sec. 2 defines the model with highly entangled GS in two directions by filling a 2D square lattice with a Motzkin chain twisted from the center. Sec. 3 analyzes the asymptotic scaling of the EE by mapping bipartitions of the 2D system into disjoint segments along the 1D chain. The result together with the scaling behavior of spectral gap converted from the 1D results for different phases and the critical points are compared with those for the previous coupled-chain constructions. Sec. 4 highlights the induced antiferromagnetic order in the spin sector along the radial direction of the 2D lattice. Sec. 5 introduces a new TN representation of the highly entangled 1D chains, the 2D twisted version of which is compared with the genuine 2D TN with similar global structure. Finally, a conclusion is given in Sec. 6 with outlook on higher dimensional generalizations and EE scaling laws for gapless GSs.
2 The twisted Motzkin chain
The $q$-deformed bicolor Motzkin chain [15] has a five-dimensional local Hilbert space, with spin up and down, both in either red (r) or blue (b) color, and a colorless spin zero. The Hamiltonian is expressed in terms of projectors
$P_{j}^{c}=|L^{c}\rangle_{j,j+1}\langle L^{c}|_{j,j+1}+|R^{c}\rangle_{j,j+1}\langle R^{c}|_{j,j+1}+|F^{c}\rangle_{j,j+1}\langle F^{c}|_{j,j+1},$ (1)
where
@@ 2 The twisted Motzkin chain
Since the anisotropic 2D model is constructed by twisting a 1D chain with a length that is quadratic in the linear size of the 2D lattice, we only need to specify the bonds on the 2D lattice occupied by degrees of freedom along the 1D chain. This is shown in Fig. 1 (a), where the colored spins reside on the lattice sites of an $9\times 8$ square lattice, with Motzkin interaction turned on along the solid edges tracing out a Motzkin chain beginning at the northwest corner and ending at the southeast corner of the lattice. Alternatively, the spins can live on the edges of the square lattice covered by the twisted Motzkin chain, and interactions are only turned on between adjacent spins on the chain, as shown in Fig. 1 (b). Depending on the location of the spin in the square lattice, it can be either between east-west (resp. north-south) neighbors for the upper and lower (resp. left and right) quadrants. Near the two diagonal lines, where the winding 1D chain turns, the interaction is between neighbors around a corner. The two layouts of the spins and neighboring interactions are asymptotically equivalent in terms of the scaling behavior of bipartite EE. However, the latter has a more regular distribution of segments along the Motzkin chain, so for convenience we will work with this lattice in the following.
@@
Figure 1: (a) A square lattice with $L(L+1)$ spins on the vertices. (b) An $L\times L$ square lattice with spins on the edges of the square lattice that are covered by a Motzkin chain of length $2N=L(L+2)$, both shown here for $L=8$. Four symmetric bipartition cuts are shown in different colors. (c) The two subsystems resulting from the yellow cut in (b). The northeast subsystem is represented with the thickened black lines, whereas the southwest is marked by thin gray lines. This is asymptotically equivalent to the yellow bipartition in (a), separating the upper half plane from the lower. The configuration depicted corresponds to the probabilistically dominant Motzkin path in the GS superposition for the $q>1$ phase, to illustrate the entanglement across subsystems. Typical height configuration at the critical point is depicted in blue. (d) The uncoupled arrays of orthogonal chains with different EE and gap scaling for comparison in table˜1 and table˜2.
@@
To study the isotropic EE, four symmetric bipartition cuts are shown in figure˜1 (a) and (b). They separate the spins into two subsystems consisting of increasingly larger segments of size 4na4n-a, for $n=1,2,\cdots,L/2+1$, with $a=0,1,2,3$ corresponding respectively to the yellow, blue, green and red cut. Starting from the center, segments within one subsystem in the 2D lattice on a given side of the cut alternate between the first and second halves of the 1D chain when it is stretched straight. Obviously the scaling behavior of bipartite EE for any of the cuts will not differ qualitatively, so from now on we focus on the yellow cut along the main diagonal with $a=0$, which results in a bipartition shown in figure˜1 (b).
3 Scaling of EE and spectral gap
@@
As analyzed in Ref. [15], the EE of the Motzkin GS decomposes into a leading contribution from the color degree of freedom, and a subleading contribution from the spin sector. Only up-down spin pairs belonging to different subsystems account for the former. In the Motzkin path picture, this means that only the net height change of a subsystem, or the surplus of up spins matters. For the 1D system studied there, the leading contribution to the EE is therefore proportional to the height at the center of the Motzkin path, giving the three distinct scaling behaviors in the two different phases and the critical point. For the critical point at $q=1$, the EE of a single connected segment in the middle of the chain has also been obtained from combinatorial analyses for the colorless [31] and colored Motzkin chains [32]. Here, the situation seems a bit more complicated, as segments belonging to two subsystems alternate along the chain. A rigorous estimation would rely on the enumeration of Motzkin-like paths that stay within the upper half plane but starts and ends at nonzero height. In Appendix A, we point out the challenge of a rigorous estimation from the combinatorial and path-integral perspective.
The decomposition of EE into a color contribution and a height fluctuation contribution can be seen from the Schmidt decomposition of the GS
[…]
This is in proportion to the the total height change in the subsystem.
Instead of evaluating the path integral equation˜A.5 by brute force, we will use the known results on the asymptotic scaling of the height function. For $q=1$, the height at distance $x$ away from the endpoint $h_{k}=\sum^{k}_{i=1}S^{z}_{i}$ on average scales as $\langle h_{k}\rangle\sim\sqrt{k}$ in the continuous limit [15, 16], which is due to the repellent hard wall at height 0. For the particular bipartition in question, we can further use the symmetry about the midpoint $\langle h_{k}\rangle=\langle h_{L-k}\rangle$ to write equation˜A.5 as
$S\sim\sum_{k=1}^{\frac{L}{2}}(-1)^{k}\langle h_{k-1}\rangle+\sum_{k=\frac{L}{2}}^{L-1}(-1)^{k}\langle h_{L-k}\rangle=\langle h_{\frac{L}{2}}\rangle=\sqrt{\sum_{j=0}^{\frac{L}{2}-1}(2L-4j)}=\sqrt{L^{2}+2L},$ (8)
which scales linearly with $L$ asymptotically. In fact, this is nothing but the average height at the middle of the chain, meaning that the bipartite EE of the 2D lattice of linear size $L$ scales in the same way as the EE of half chain with length $L^{2}$.
For the $q>1$ phase, $\langle h_{k}\rangle\sim k$, so a completely analogous argument gives the scaling $S\sim L^{2}$. The $q<1$ phase, on the other hand, has a bounded height everywhere $\langle h_{k}\rangle\sim\mathcal{O}(1)$, which results in the scaling $S\sim\mathcal{O}(1)$. Notice that this EE is significantly sub-area law, reminiscent of the colorless version of the coupled chain models [22, 21], and more severe than the topological and logarithmic corrections of the colored loop models studied in Ref. [20].
To estimate the EE scaling in the various phases of the Motzkin GS we can also use its matrix product state (MPS) representation, found in Ref. [28]. By summing up the absolute value of the height differences s across segments $s$ in one of the subsystems defined by one of the cuts in figure˜1, we can estimate the EE $S$. That is
$S\sim\sum_{s\in{\mathcal{S}_{\text{cut}}}}|{}_{s}|=\sum_{s\in{\mathcal{S}_{\text{cut}}}}|\langle h^{s}_{f}-h^{s}_{i}\rangle|,$ (9)
where $\mathcal{S}_{\text{cut}}$ contains all segments in one of subsystems defined by one of the colored cuts in figure˜1 (a) and $h^{s}_{i}$ (resp. $h^{s}_{f}$) is the height at the beginning (resp. end) of the segment $s$. For example, if we consider the blue cut in figure˜1 (a) and the northeast subsystem, $\mathcal{S}_{\text{cut}}$ will consist of segments $1\dots 15]$, $47\dots 57]$, $27\dots 33]$ and $[37,38,39]$. In each of the segments, $h^{s}_{i}$ and $h^{s}_{f}$ is the height at the spin denoted by the first and the last number respectively. Note that equation˜9 is identical with the first form of equation˜A.5. The expected height values $\langle h\rangle$ are computed from the MPS representation. This gives the plots in figure˜2. Clearly, the scaling in the various phases matches the above results and are independent on the exact bipartition considered in the large $L$ limit.
Figure 2: Scaling of the bipartite EE $S$ computed from equation˜9 and the MPS representation of the GS for various linear system sizes $L$ and for the different colored cuts in figure˜1 (a). In (a) $q=0.99$, in (b) $q=1$ and in (c) $q=1.01$.
@@
The subleading contribution to the EE is from the fluctuation in the height of Motzkin paths, which is also the sole source of entanglement for a monochromatic version of the model. A Schmidt decomposition of the reduced density matrix shows that the two subsystems only depend on each other via the total spin within one subsystem, such that the total spin of the whole chain is zero. Moving from the 1D system to our 2D construction, nothing changes except the subsystem is distributed into disconnected segments. So the dimensionality of the reduced density matrix is still the number of possible values of the total spin. However, in the $q>1$ (resp. $q<1$) phase, the probability of the total spin, or height change decays exponentially with the deviation of the area underneath a Motzkin path from the most probable one that is of highest height (resp. completely flat) [15, 16]. So only paths with a height change at finite distance away from the typical one effectively matters, giving a contribution to the EE from height fluctuations that is a bounded constant independent of system size for both phases. When $q=1$, there is more fluctuation. Although entropic repulsion pushes the average height to order $\sqrt{N}$ above the hard wall at zero height, the variance of the Gaussian free field remains to scale linearly with the total length of the segments, making the EE scale as $\log N$ [11]. Note that in a monochromatic version of the 2D model, all three scalings are sub-area law, which is unique to our construction of filling the 2D lattice with a twisted 1D chain.
It is worth comparing the EE scalings of the two different 2D models together with a quasi-1D model composed of two layers of Motzkin chains in orthogonal directions that are not interacting with each other, as shown in figure˜1 (d). Such a construction correspond to yet another way to define height function in 2D, namely by the Cartesian product of heights along the chains in two different directions. Since the chains in each layer are neither interacting with one another, nor with those in the other direction, the EE scaling in both phase and the at the critical point are simply the corresponding scaling of a single 1D chain multiplied by $L$, as summarized in table˜1.
@@
Table 1: Comparison of the EE scalings with linear system size $L$ between the two different 2D constructions and the quasi-1D arrays of uncoupled chains.
Model Phase Coupled chains Twisted chain Uncoupled chains
$q>1$ $L^{2}$ [21, 22] $L^{2}$ $L^{2}$
Colored $q=1$ $L\log L$ [21, 22] $L$ $L^{\frac{3}{2}}$
$q<1$ $L$ [21, 22] $\mathcal{O}(1)$ $L$
$q>1$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ $L$
@@
Colorless $q=1$ $L$  [^1] This is an upper-bound from the zeroth order Rényi entropy. The random membrane when bipartitioned gives a Motzkin path on the boundary, giving the dimensionality of the reduced density matrix. The same upper-bound was also obtained for six-vertex models with domain wall boundary conditions using the exact enumeration of alternating sign matrices [33]. $\log L$ $L\log L$
$q<1$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ $L$
@@
Unlike the scaling of EE discussed above, which requires somewhat different analyses than the 1D system due to the more complicated bipartition, upper-bounds on the scaling of the spectral gap can be straightforwardly deduced from the 1D scaling simply by substituting the linear system size with $N=L(L+2)$. The results are summarized in table˜2. This should be compared with the EE scaling in table˜1. In particular, the $q>1$ phase of the colored model shows that the same EE scaling does not imply the same scaling of spectral gap, and vice versa as shown by the critical point of the colored model.
Table 2: Comparison of the spectral gap scalings with linear system size $L$ between the two different 2D constructions and the quasi-1D arrays of uncouples chains, where asterisks denote conjecture scalings. The greek letters denote different positive constants.
Model Phase Coupled chains Twisted chain Uncoupled chains
[…]
$q<1$ $\mathcal{O}(1)*$ [30] $\mathcal{O}(1)$ [35] $\mathcal{O}(1)$ [35]
[…]
$q<1$ $\mathcal{O}(1)*$ [30] $\mathcal{O}(1)$  [35] $\mathcal{O}(1)$  [35]
4 Antiferromagnetic order in the radial direction